Mar 19,2013 Readout

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST

  • ESNET (Number is 8542553) and EVO session (GlueX Calorimetry meeting room)
  • Phone connection only upon request.
    • +1-866-740-1260 : US and Canada
    • +1-303-248-0285 : International
    • then use participant code: 3421244# (the # is needed when using the phone)
    • OR from type access code 3421244 into "join a meeting" without the # (you need java plugin)

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Videoconference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Items for followup from previous meeting

Presentations from recent meetings

Tentative Agenda

  1. Announcements
  2. Review of Action Items
  3. Monitoring
    • Note from Christina: I am in Portugal-but there is nothing to report on BCAL.The FCAL boards will start getting stuffed probably by next week
  4. Reconstruction and simulation
  5. Update from JLab
  6. Update from USM
  7. Status of testing at Regina
  8. Beam Tests, Cosmics
  9. Other


Attending:Elton, Tim, Shaun, Yi, David, Beni (Jlab); Andrei (UofR); Hayk, Alam, Orlando, Elias (USM).

  1. Announcements
  2. Review of Action Items
  3. Monitoring
    • Note from Christina: I am in Portugal-but there is nothing to report on BCAL.The FCAL boards will start getting stuffed probably by next week
  4. Reconstruction and simulation
    1. Beni (first comment)
      • Had some questions so started looking at the reconstruction
      • Parameters in HDGEANT, which are also neede din reconstruction (e.g. attenuation length, veff, etc) are numbers in the code. Other systems have moved these types of parameters to the data base. The Bcal should also do this
    2. Andrei
      • The main issue is clarifying and deciding on the algorithm that is to be used
      • Previous code used unrealistic values for simulation
      • Now implementing amplitude smearing in a more realistic way
      • Existing code does not correctly identify clusters
      • Suggests using physics shapes to use as patterns in the cluster recognition code. Problem is to distinguish between multiple clusters and background
    3. Beni (second comment)
      • HDGEANT applies a threshold to the data before MC smear. In principle, one should take thresholds after the smearing has been applied.
      • Andrei: These thresholds are very low and it is very unlikely that any eliminated pulse would survive to the detector
    4. David
      • Meetings on Thursdays continue to discuss the reconstruction.
      • Will is looking at time-walk corrections using a self-developed algorithm. He is looking into an another algorithm proposed by Andrei.
      • Another problem found in the data are the single p.e. peaks visible in the energy spectrum, which should be washed out.
  5. Update from JLab
    • Production gluing continues at a rate of 4modules/week.
    • A stack of boxes with completed modules are in the hall and they are beginning the process of placing them on to the cradle
    • Module 45 (second one glued) was taken out of the box and found to have dark spots on the interface between the guide and the module if viewed at a particular angle. The module was taken back to the ESB, where we are doing the gluing for inspection. It dark spots appeared on one end of the module on several guides. The guides were removed and re-glued. The other side did not have similar signs. The module will be re-tested.
    • Other modules will be inspected carefully as they are taken out of their boxes.
  6. Update from USM
    • File:Status Report 20130319.pdf
    • Testing of final MPPCs: to date 305 have been tested successfully, and 65 are under test.
    • Expect to complete testing this week
    • A box with 640 MPPCs is ready to ship now
    • The final box with 370 MPPCs will be ready to ship next week.
    • Three MPPCs did not meet specifications. Hayk has contacted Hamamatsu for replacement. They have sent him RMA paperwork to complete. Hayk and Elton will work offline to expedite replacement. There is a question about how to complete the shipment of 370 units next week. Elton suggested one possibility would be to include the non-conforming MPPCs and JLab could make the exchange with Hamamatsu, but this would have to be worked out.
  7. Inventory of SiPMs
    • Media: SiPM serial numbers.png
    • JLab also needs the MPPCs sent out on loan from the first-article batch. USM has 40 and U of R has 10.
    • Andrei : When does JLab need them? Elton: At the time that the other large shipments are completed, in about 2 weeks.
    • Andrei asked if they could keep them for 3-4 weeks. Elton: This might start having an impact on schedule, but it depends on the details for sorting and completing the readout assemblies.
  8. Status of testing at Regina
    • The test setup was found to have an asymmetric response. The problem was traced to a kinked fiber on one side of the detector. SIPMs are being remeasured to qualify the test setup.
  9. Beam Tests, Cosmics
  10. Other